OpenStudy (anonymous):

PLEASE HELP WILL FAN AND MEDAL below is the graph of f(x)=3log(base10)x how would you describe the graph of g(x)=9log(base10)x? Ag(x) compresses f(x) by a factor of 3 B g(x) shifts f(x) up 3 units C g(x) shifts f(x) to the right 3 units D g(x) stretches f(x)vertically by a factor of 3

OpenStudy (tkhunny):

No shifting.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

we can write this: \[g(x)=9\log _{10}x=3(3\log _{10}x)=3f(x)\] so, in order to draw the graph of g(x), it is necessary multiply by 3 the y-coordinate of all points of the graph of f(x)

OpenStudy (michele_laino):

so D.

OpenStudy (emmatassone):

D.

OpenStudy (anonymous):

thank you @Michele_Laino

OpenStudy (michele_laino):

thank you @Dianie